OpenStudy (anonymous):

does anyone know how to solve this: 3^x+3=1/9 step by step and not just the answer

OpenStudy (kropot72):

\[\large 3^{(x+3)}=\frac{1}{9}\] Please confirm that this is the question.

OpenStudy (anonymous):

yes, that is the question : )

OpenStudy (kropot72):

\[\large3^{-1}=\frac{1}{3}\] \[\large3^{-2}=\frac{1}{3}\times\frac{1}{3}=\frac{1}{9}\] Can you see that: x + 3 = -2 ............(1) Now can you find the value of x by solving equation (1)?

OpenStudy (anonymous):

okay, I got it thanks : ))

OpenStudy (mathmath333):